### v2

@@ 3.2.1 Number of rainy days/rain day frequency
The line plots in Figures 6 and 7show the MEs on the number of rainy days of each product at the various stations in Zambia and Ghana respectively. The height of the bars at the different stations represents the mean number of rainy days at those stations. Each bar is divided into four equal segments (separated by horizontal white lines), with the first segment at the bottom and the fourth segment at the top. These divisions facilitate the estimation of PBIAS. For example, TAMSAT has an ME of about 40 mm at Chipata while having a PBIAS of about 50% (see Figure 6).
@@
On average, all the REs overestimated the number of rainy days in Zambia and Ghana (Figures 6 and 7).
@@
Figures 8 and 9 illustrate the performs of the REs in detecting rainy days across various stations in Zambia and Ghana respectively.
@@
From Figures 8 and 9, it was observed that all the products had considerable proportions of False Positives. This occurrence was lowest in ENACTS, followed by CHIRPS and TAMSAT. Proportion of False Negatives were lowest in ERA5, AGERA5, and PCDR and was relatively higher in CHIRPS, ENACTS, and CHIRP.
@@ 3.2.2 Annual rainfall totals
As compared to rainfall frequency, the biases in total rainfall were generally low for all the REs both in Zambia and Ghana ($-30$% $\leq$ PBIAS $\leq$ 30%) as can be seen in Figures 10 and 11 respectively. CHIRPS, TAMSAT and CHIRP had the least biases ($-20$% $<$ PBIAS $\leq$ 10%) in Zambia. ENACTS, CHIRPS, CHIRP, and TAMSAT underestimated total annual rainfall at most of the stations in Zambia while ERA5 and AGERA5, overestimated total rainfall at most of the stations in Zambia (Figure 10). PCDR overestimated at some of the stations (0% $<$ PBIAS $<$ 20%) while underestimating at some (with $-20$% $<$ PBIAS $<$ 0%). The product with the highest underestimation was ENACTS ($-30$% $\leq$ PBIAS $<$ 0%) across all stations while AGERA5 was the product with the highest overestimation (0% $<$ PBIAS $\leq$ 30%) at most of the stations.
@@
The REs tended to have less variability than the observed rainfall total (RSD $<$ 1) across most of the stations in Zambia and Ghana (Figures 10 and 11 respectively).
@@ 3.2.3 Mean rain per rainy day
The REs underestimated mean rain per rainy day at varying degrees, with $-60$% $<$ PBIAS $<$ 0% at most of the stations in Zambia (Figure 12), and $-75$% $<$ PBIAS $<$ 0 at most of the stations in Ghana (Figure 13). Correlations were generally low for most of the products across all the stations. Aside from ENACTS (with 0.5 $<$ r $<$ 1 at most of the locations), most of the REs had correlations between $-0.3$ and 0.5 at most of the stations in Zambia and Ghana (respectively Figures 12 and 13). The REs also tended to have lesser variability (RSD $<$ 0.8) across most of the stations as compared to gauge observations in Zambia and Ghana.
@@ 3.3 Comparison of Seasonal Behaviour
Figures 14-20 show the Zero-order Markov chain models with three harmonics fitted separately to the rainfall occurrence at each of the stations in Zambia and Ghana for the gauge observations as well as the REs. At each station, only days which the gauge observations and RE values were not missing were included to ensure comparison of the same days. The solid black curve shows the fitted rain day frequency for the gauge observations with 0.85mm rain day threshold. The coloured dashed lines show the fitted rain day frequency for the RE values at various rain day thresholds (0.85mm, 2mm, 3mm, 4mm, and 5mm). The y-axis on each plot represents the proportion of rain day frequency, while the x-axis shows the date in the year, spanning from August 1 to July 31 for Zambia, and January 1 to December 31 for Ghana.
Figure 14: Rain day frequency models of gauge observations vs ERA5 in Zambia
@@
The seasonality at the Zambian stations were well captured by all the REs, with the highest rain day frequency in the November to May season (see Figures 14-17 for ERA5, TAMSAT, CHIRPS, and ENACTS respectively consistent with the REs not shown here).
@@
Aside from CHIRPS and PCDR, the REs also captured seasonality of the northern stations of Ghana well, capturing the unimodal patterns. CHIRPS and PCDR (Figures 20 and 18 respectively) exhibited bi-modal patterns at these northern stations, which have unimodal rainfall patterns.
@@ 3.4 Rainfall Intensity
Figure 21 and 22 show the different REs and their probability of detection of different rainfall intensity categories at the different stations in Zambia and Ghana respectively.
Figure 21: Probability of detection for different rainfall intensity categories in Zambia
Figure 22: Probability of detection for different rainfall intensity categories in Ghana
All the REs showed high skill of detecting dry days (with 70% $<$ POD $<$ 100%) across all stations in Zambia (Figure 21). ENACTS, CHIRPS and TAMSAT had about 10% chance higher in detecting dry days than the other REs. Apart from CHIRP, all the REs had POD $>$ 50% in detecting moderate rain at almost all the stations, with ENACTS having 50% $<$ POD $<$ 80%, outperforming all the REs, followed by TAMSAT (with 50% $<$ POD $<$ 70%). All the REs had a probability of less than 50% in detecting light rain. They all performed poorly in detecting heavy and violent rains (with POD close to zero across all stations).
In the case of Ghana, not all the REs demonstrated a high probability of accurately detecting dry days, as shown in Figure 22. Aside from TAMSAT and CHIRPS, which exhibited a POD of at least 70% across all stations, the other REs had POD often below 70% in southern Ghana, and 60% $<$ POD $<$ 85% in northern Ghana. The two reanalysis products — ERA5 and AGERA5 — stood out slightly in the detection of light rain, even though they both had a POD below 50% at most of the stations. TAMSAT outperformed all the REs in the detection of moderate rain (with 50% $<$ POD $<$ 80%) at almost all the stations. The worst performance of most of the REs in detecting dry days was observed in stations located in the central and southern regions, particularly along the southwestern coast of the country.
Similar to Zambia, the REs performed poorly in the detection of heavy and violent rains (with POD close to zero across all stations).
Figure 23: Graph of Kasama station rainfall intensity categories
@@
Figure 23 represents the distribution of hits and misses in the different rainfall intensity categories at Kasama, Zambia. True Hit represents the proportion of rainfall events in the rainfall intensity category that were correctly detected by the RE; True Miss represents the proportion of rainfall events in the rainfall intensity category that the RE detected as a dry day. Lower represents the proportion of rainfall events in the rainfall intensity category that the RE missed by estimating a lower intensity category. Higher represents the proportion of rainfall events in the rainfall intensity category that the RE missed by estimating a higher intensity category. The distribution of hits and misses in Kasama was not very different from those at other stations in Zambia and Ghana (not shown). True Hit was relatively high for moderate rain, which relates to the observation in Figure 21. As expected, True Miss was very low across all intensity categories. True Miss was the lowest under violent rain followed by heavy rain. The figure further shows that rainfall was correctly detected by most of the REs on days of violent and heavy rain, but the intensity was usually rather underestimated.
@@ 4 Discussion
    Can the product reliably monitor extreme rainfall events in region X?
@@
This targeted validation enables more informed and context-specific decision-making.
The topography in the southern part of Ghana (including locations like Axim and Takoradi) is complex due to the vegetation and terrain, and its proximity to the coast and surrounding forest regions. These regions are subject to localized weather patterns [3], which can be harder for satellite-based or reanalysis products to accurately capture. Coastal influences, such as sea breezes, can induce small-scale convective systems that could be missed by the REs. On the contrary, Zambia is predominantly flat and landlocked, with rainfall largely driven by convective systems, and there is typically a strong correlation between the cold cloud tops of these systems and rainfall [42]. Our results confirm the fact performance of the REs under annual summaries, seasonality studies, and rainfall intensity detection was more consistent in Zambia as compared to Ghana.
Previous works have shown that REs are usually good at detecting dry days at many places [46, 64]. From our results, all the REs were seen to detect dry days well in Zambia and northern Ghana while CHIRPS and TAMSAT still maintained consistent high performance in south Ghana as well, consistent with previous existing literature. One explanation could be as explained in the previous paragraph. Another reason could be due the fact that TAMSAT and CHIRPS have been designed for drought monitoring [22, 42].
It is known that the REs are extensively validated on their ability to capture daily, sub-seasonal, seasonal, or annual rainfall amounts during development [22, 42, 4]. Our results shown on Figures 10 and 11 show low biases and good correlations on total rainfall, consistent with [2], and likely explained by the literature above.
@@
ENACTS includes a large number of meteorological station data than other similar blended REs [56, 18]. CHIRPS also includes some station data, while TAMSAT is calibrated with station data. Our results showed that ENACTS stood out as one of the best products across multiple contexts in Zambia (followed by CHIRPS and TAMSAT). This is likely due to the merging/calibration of these products with station data. Local calibration could be a potential approach for reducing the satellite precipitation errors, as was also noted by [14] and [15].
REs have been known with their tendency to overestimate the number of rainy days, as confirmed by previous works within the Africa region [44]. Our results as shown in Figures 6 and 7 is consistent with existing findings. This tendency to overestimate rainy days is expected, as the REs represent area-based precipitation estimates, while gauge observations are point measurements. Due to the overestimation, they had low instances of False Negatives. This is confirmed by the results in Figures 8 and 9 for Zambia and Ghana respectively. They also had considerable proportions of False Positives due to the overestimation (see Figures 8 and 9). Some of these False Positives may have stemmed from rainfall occurring elsewhere within the grid area but outside the immediate vicinity of the gauge. Further investigation, using additional station data from within these grids, could clarify these discrepancies. It is also possible that rainfall detected by REs occurs several kilometers away or even on a different day, as noted by [32]. Rain-bearing clouds can travel hundreds of kilometers before releasing rain, sometimes accumulating over several days before delivering intense rainfall in a single event [27].
All the REs showed poor performance in detecting heavy and violent rainfalls (see Figures 21 and 22), which is consistent with existing literature [62, 46, 64, 2]. This indicates that the products in their current state are likely unsuitable for extreme rainfall detection or studies at these locations. Developing products tailored to extreme rainfall detection, which could then complement existing products, is imperative. [20] noted that these extreme events have not been extensively explored. As illustrated in Figure 23, most heavy and violent rain days were correctly classified as rainy by the products, albeit with underestimated rainfall amounts. This suggests that focused enhancements in design could significantly improve detection accuracy and rainfall estimation for these events.
@@ 5 Conclusion
This study evaluated the performance of eight satellite and reanalysis rainfall products — CHIRPS, TAMSAT, CHIRP, ENACTS, ERA5, AgERA5, PERSIANN-CDR, and PERSIANN-CCS-CDR — in Zambia and Ghana using a point-to-pixel validation approach. The analysis encompassed spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection across 38 ground stations. The results have provided useful insights into the performance of the products and the implication on their usage as follows:
All products exhibited high probability of detection (POD) for dry days in Zambia and nothern Ghana (with 70% $<$ POD $<$ 100%, and 60% $<$ POD $<$ 85% respectively) while CHIRPS and TAMSAT maintained a consistent high performance in southern Ghana. These products may potentially be useful for drought detection at those locations.
There were biases in the products under multiple summaries (including total rainfall, rain day frequency). Bias-correction is recommended, especially for reducing the biases on the number of rainy days.
@@
Lastly, all REs performed poorly in the detection of heavy and violent rainfall (with POD close to 0%). At their current state, the products may not be recommended for extreme rainfall events detection at these locations, such as floods. Future research should look more into this.
@@
Not applicable

### v4

@@ 3.2.1 Number of rainy days/rain day frequency
The line plots in Figures 8 and 9show the MEs on the number of rainy days of each product at the various stations in Zambia and Ghana respectively. The height of the bars at the different stations represents the mean number of rainy days at those stations. Each bar is divided into four equal segments (separated by horizontal white lines), with the first segment at the bottom and the fourth segment at the top. These divisions facilitate the estimation of PBIAS. For example, TAMSAT has an ME of about 40 mm at Chipata while having a PBIAS of about 50% (see Figure 8).
@@
On average, all the REs overestimated the number of rainy days in Zambia and Ghana (Figures 8 and 9).
@@
Figures 10 and 11 illustrate the performs of the REs in detecting rainy days across various stations in Zambia and Ghana respectively.
@@
From Figures 10 and 11, it was observed that all the products had considerable proportions of False Positives. This occurrence was lowest in ENACTS, followed by CHIRPS and TAMSAT. Proportion of False Negatives were lowest in ERA5, AGERA5, and PCDR and was relatively higher in CHIRPS, ENACTS, and CHIRP.
@@ 3.2.2 Annual rainfall totals
As compared to rainfall frequency, the biases in total rainfall were generally low for all the REs both in Zambia and Ghana ($-30$% $\leq$ PBIAS $\leq$ 30%) as can be seen in Figures 12 and 13 respectively. CHIRPS, TAMSAT and CHIRP had the least biases ($-20$% $<$ PBIAS $\leq$ 10%) in Zambia. ENACTS, CHIRPS, CHIRP, and TAMSAT underestimated total annual rainfall at most of the stations in Zambia while ERA5 and AGERA5, overestimated total rainfall at most of the stations in Zambia (Figure 12). PCDR overestimated at some of the stations (0% $<$ PBIAS $<$ 20%) while underestimating at some (with $-20$% $<$ PBIAS $<$ 0%). The product with the highest underestimation was ENACTS ($-30$% $\leq$ PBIAS $<$ 0%) across all stations while AGERA5 was the product with the highest overestimation (0% $<$ PBIAS $\leq$ 30%) at most of the stations.
@@
The REs tended to have less variability than the observed rainfall total (RSD $<$ 1) across most of the stations in Zambia and Ghana (Figures 12 and 13 respectively).
@@ 3.2.3 Mean rain per rainy day
The REs underestimated mean rain per rainy day at varying degrees, with $-60$% $<$ PBIAS $<$ 0% at most of the stations in Zambia (Figure 14), and $-75$% $<$ PBIAS $<$ 0 at most of the stations in Ghana (Figure 15). Correlations were generally low for most of the products across all the stations. Aside from ENACTS (with 0.5 $<$ r $<$ 1 at most of the locations), most of the REs had correlations between $-0.3$ and 0.5 at most of the stations in Zambia and Ghana (respectively Figures 14 and 15). The REs also tended to have lesser variability (RSD $<$ 0.8) across most of the stations as compared to gauge observations in Zambia and Ghana.
@@ 3.3 Comparison of Seasonal Behaviour
Figures 16-22 show the Zero-order Markov chain models with three harmonics fitted separately to the rainfall occurrence at each of the stations in Zambia and Ghana for the gauge observations as well as the REs.
Figure 16: Rain day frequency models of gauge observations vs ERA5 in Zambia. At each station, only days which the gauge observations and RE values were not missing were included to ensure comparison of the same days. The solid black curve shows the fitted rain day frequency for the gauge observations with 0.85mm rain day threshold. The coloured dashed lines show the fitted rain day frequency for the RE values at various rain day thresholds (0.85mm, 2mm, 3mm, 4mm, and 5mm). The y-axis on each plot represents the proportion of rain day frequency, while the x-axis shows the date in the year, spanning from August 1 to July 31. Figures 17-22 were created in a similar way except that a year at the Ghanaian stations start from January 1 and end on December 31
@@
The seasonality at the Zambian stations were well captured by all the REs, with the highest rain day frequency in the November to May season (see Figures 16-19 for ERA5, TAMSAT, CHIRPS, and ENACTS respectively consistent with the REs not shown here).
@@
Aside from CHIRPS and PCDR, the REs also captured seasonality of the northern stations of Ghana well, capturing the unimodal patterns. CHIRPS and PCDR (Figures 22 and 20 respectively) exhibited bi-modal patterns at these northern stations, which have unimodal rainfall patterns.
@@ 3.4 Rainfall Intensity
Figure 23 and 24 show the different REs and their POD for different rainfall intensity categories at the different stations in Zambia and Ghana respectively.
Figure 23: POD for different rainfall intensity categories in Zambia
Figure 24: POD for different rainfall intensity categories in Ghana
3.4.1 Dry days detection
All the REs showed high skill of detecting dry days (with 70% $<$ POD $<$ 100%) across all stations in Zambia (Figure 23). ENACTS, CHIRPS and TAMSAT had about 10% chance higher in detecting dry days than the other REs. In the case of Ghana, not all the REs demonstrated a high probability of accurately detecting dry days, as shown in Figure 24. Aside from TAMSAT and CHIRPS, which exhibited a POD of at least 70% across all stations, the other REs had POD often below 70% in southern Ghana, and 60% $<$ POD $<$ 85% in northern Ghana. The worst performance of most of the REs in detecting dry days was observed in stations located in the central and southern regions, particularly along the southwestern coast of the country.
3.4.2 Light and moderate rains detection
In the case of Zambia all the REs (with the exception of CHIRP) had POD $>$ 50% in detecting moderate rain at almost all the stations, with ENACTS having 50% $<$ POD $<$ 80%, outperforming all the REs, followed by TAMSAT (with 50% $<$ POD $<$ 70%). All the REs had a probability of less than 50% in detecting light rain.
In the case of Ghana, the two reanalysis products — ERA5 and AGERA5 — stood out slightly in the detection of light rain, even though they both had a POD below 50% at most of the stations. TAMSAT outperformed all the REs in the detection of moderate rain (with 50% $<$ POD $<$ 80%) at almost all the stations.
3.4.3 Heavy and violent rains detection
All the REs all performed poorly in detecting heavy and violent rains (with POD close to zero across all stations) in both Ghana and Zambia (see Figure 23 and 24).
Figure 25: Graph of Kasama station rainfall intensity categories. True Hit represents the proportion of rainfall events in the rainfall intensity category that were correctly detected by the RE; True Miss represents the proportion of rainfall events in the rainfall intensity category that the RE detected as a dry day. Lower represents the proportion of rainfall events in the rainfall intensity category that the RE missed by estimating a lower intensity category. Higher represents the proportion of rainfall events in the rainfall intensity category that the RE missed by estimating a higher intensity category.
@@ 3.4.3 Heavy and violent rains detection
Figure 25 represents the distribution of hits and misses in the different rainfall intensity categories at Kasama, Zambia. The distribution of hits and misses in Kasama was not very different from those at other stations in Zambia and Ghana (not shown). True Hit was relatively high for moderate rain, which relates to the observation in Figure 23. As expected, True Miss was very low across all intensity categories. True Miss was the lowest under violent rains followed by heavy rain. The figure further shows that rainfall was correctly detected by most of the REs on days of violent and heavy rains, but the intensity was usually rather underestimated.
@@
These results highlight the strengths and limitations of REs, with implications for their application in climate studies such as drought, and extreme event analysis. The results are discussed in the next section.
@@ 4 Discussion
    Can the product reliably monitor heavy and violent rain events in region X?
@@
This targeted validation enables more informed and context-specific decision-making. For instance, improved validation of REs could help farmers optimize irrigation strategies, reduce crop losses due to drought.
The topography in the southern part of Ghana (including locations like Axim and Takoradi) is complex due to the vegetation and terrain, and its proximity to the coast and surrounding forest regions. These regions are subject to localized weather patterns [3], which can be harder for the REs to accurately capture. Coastal influences, such as sea breezes, can induce small-scale convective systems that could be missed by the REs. On the contrary, Zambia is predominantly flat and landlocked, with rainfall largely driven by convective systems, and there is typically a strong correlation between the cold cloud tops of these systems and rainfall [47]. Our results confirm the fact performance of the REs under annual summaries, seasonality studies, and rainfall intensity detection was more consistent in Zambia as compared to Ghana.
Previous works have shown that REs are usually good at detecting dry days at many places [51, 70]. From our results, all the REs were seen to detect dry days well in Zambia and northern Ghana while CHIRPS and TAMSAT still maintained consistent high performance in south Ghana as well, consistent with previous existing literature. One explanation could be as explained in the previous paragraph. A reason for the superior performance of CHIRPS and TAMSAT on detecting dry days could be due the fact that they have been designed for drought monitoring [25, 47].
It is known that the REs are extensively validated on their ability to capture daily, sub-seasonal, seasonal, or annual rainfall amounts during development [25, 47, 4]. Our results shown on Figures 12 and 13 show low biases and good correlations on total rainfall, consistent with [2], and likely explained by the literature above.
@@
ENACTS includes a large number of meteorological station data than other similar blended REs [62, 20]. CHIRPS also includes some station data, while TAMSAT is calibrated with station data. Our results showed that ENACTS stood out as one of the best products across multiple contexts in Zambia (followed by CHIRPS and TAMSAT). This is likely due to the merging/calibration of these products with station data. Local calibration could be a potential approach for reducing the satellite precipitation errors [7, 16, 17].
REs have been known with their tendency to overestimate the number of rainy days, as confirmed by previous works within the Africa region [49]. Our results as shown in Figures 8 and 9 are consistent with existing findings. This tendency to overestimate rainy days is expected, as the REs represent area-based precipitation estimates, while gauge observations are point measurements. Due to the overestimation, they had low instances of False Negatives. This is confirmed by the results in Figures 10 and 11 for Zambia and Ghana respectively. They also had considerable proportions of False Positives due to the overestimation (see Figures 10 and 11). Some of these False Positives may have stemmed from rainfall occurring elsewhere within the grid area but outside the immediate vicinity of the gauge. Further investigation, using additional station data from within these grids, could clarify these discrepancies. It is also possible that rainfall detected by REs occurs several kilometers away or even on a different day, as noted by [36]. Rain-bearing clouds can travel hundreds of kilometers before releasing rain, sometimes accumulating over several days before delivering intense rainfall in a single event [31].
All REs exhibited poor performance in detecting heavy and violent rains (see Figures 23 and 24), consistent with findings reported in existing literature [68, 51, 70, 2]. This suggests that, in their current state, these products are likely unsuitable for detecting heavy and violent rains or for conducting related studies in these regions. The challenges in accurately capturing these events may be attributed to spatial resolution constraints, limitations in rainfall retrieval algorithms, and regional variability [7, 51]. Our results are based on a point-to-pixel validation approach, where spatially averaged RE values from their native pixels are directly compared with point-based station observations [50]. This approach may dilute the intensity of localized heavy and violent rains, leading to underestimation. Furthermore, the rainfall variability in our study regions, which is predominantly driven by convective systems, poses additional challenges for REs in detecting the heavy and violent rains associated with these systems [7].
Despite the REs’ limitation in detecting heavy and violent rains, Figure 25 demonstrates that most days of heavy and violent rains were correctly classified as rainy by the REs, albeit with underestimated rainfall amounts. This indicates that the products have the potential to detect heavy and violent rains but require targeted improvements, such as region-specific calibration, to better capture the intensity and distribution of heavy rainfall [23]. Developing REs specifically tailored to heavy and violent rains detection, which could complement existing products, is essential for enhancing climate services in regions vulnerable to these events.
@@ 5 Conclusion
This study evaluated the performance of eight REs — CHIRPS, TAMSAT, CHIRP, ENACTS, ERA5, AgERA5, PERSIANN-CDR, and PERSIANN-CCS-CDR — in Zambia and Ghana using a point-to-pixel validation approach. The analysis encompassed spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection across 38 ground stations. The results have provided useful insights into the performance of the products and the implication on their usage as follows:
All products exhibited high POD for dry days in Zambia and nothern Ghana (with 70% $<$ POD $<$ 100%, and 60% $<$ POD $<$ 85% respectively) while CHIRPS and TAMSAT maintained a consistent high performance in southern Ghana. These products may potentially be useful for drought detection at those locations.
There were biases in the products under multiple summaries (including total rainfall, rain day frequency). Bias-correction is recommended, especially for reducing the biases on the number of rainy days. Numerous studies have demonstrated effective bias-correction techniques for REs (e.g., [30, 61]), which could be adapted for this purpose. However, the application of bias-correction methods and the recommendation of specific techniques for our study regions were beyond the scope of this study.
@@
The results from this study are likely applicable to regions with similar climatic characteristics, particularly in sub-Saharan Africa and other areas reliant on rain-fed agriculture. However, we recommend validation in such locations according to the use case. The methods used in this work can be readily applied to new locations, even at one station. This makes this approach particularly valuable for regions facing similar climatic and agricultural challenges, where reliable rainfall data are critical for climate services and decision-making.
Lastly, all REs performed poorly in the detection of heavy and violent rains (with POD close to 0%). At their current state, the products may not be recommended for heavy and violent rains detection at these locations, such as floods. Future research should look more into this.
@@
Additional figures and tables are in the Supplementary Material (Tables E1-E6, and Figures E1-E69).